# Generalized convolutional many body distribution functional representations

### v1

@@ Abstract
Modern machine learning (ML) models of chemical and materials systems with billions of parameters require vast training datasets and considerable computational efforts. Lightweight kernel or decision tree based methods, however, can be rapidly trained, leading to a considerably lower carbon footprint. We introduce generalized many-body distribution functionals as highly compute and data-efficient atomic representations for accurate kernels that excel in low-data regimes. Generalizing the MBDF framework, cMBDF encodes local chemical environments in a compact fashion using translationally and rotationally invariant functionals of smooth atom centered Gaussian electron density proxy distributions weighted by interaction potentials. The functional values can be efficiently evaluated by expressing them in terms of convolutions which are calculated via fast Fourier transforms and stored on pre-defined grids. In the generalized form each atomic environment is described using a set of functionals uniformly defined by three integers; many-body-, distribution derivative- and weighting function orders. Irrespective of system size and composition, cMBDF atomic feature vectors remain compact and constant in size for a fixed choice of these orders controlling the structural and compositional resolution. While being up to two orders of magnitude more compact than other popular representations, cMBDF is shown to be more accurate for the learning of various quantum properties such as energies, dipole moments, homo-lumo gaps, heat-capacity, polarizability, optimal exact-exchange admixtures and basis-set scaling factors. Applicability for organic and inorganic chemistry is tested as represented by the QM7b, QM9 and VQM24 data sets. Due to its compactness, model training and testing times are reduced from 23 hours to 8 minutes, implying a corresponding reduction in carbon footprint. The versatility, accuracy, and computational efficiency obtained suggest that cMBDF holds great promise as a crucial ingredient for foundational yet green ML models in the chemical and materials sciences.
@@ I Introduction
For instance, carbon dioxide emissions from training large language models, for instance, are soon expected to match the monthly emissions of New York City 8.
@@ IV Conclusions
In this work we have introduced a systematically improvable class of compact atomic representations for use throughout chemical compound space. Convolutional many-body distribution functionals (cMBDF) encode the chemical environment of an atom through a set of translationally and rotationally invariant functionals of the atomic density. The functional values are efficiently evaluated via fast Fourier transforms using the convolution theorem which can be evaluated and stored on a pre-defined grid. Weighting functions of various types can be incorporated to capture physical interactions (short and long range) of interest efficiently via the atomic density functionals. The atomic feature vector size is invariant to the cut-offs employed, system size and composition leading to a compact and constant size atomic descriptor. cMBDF is shown to outperform other commonly used representations for the learning of a variety of physical properties across chemical space while remaining nearly two orders of magnitude more compact. Due to its computational efficiency and versatility, cMBDF can lead to significantly faster prototyping, training and testing of quantum machine learning models for a variety of tasks across chemical compound space. Future work will include a study of cMBDF gradients and their applicability across geometrical changes for relaxing geometries and identifying transition states 44.
@@ Acknowledgements.
D.K. is thankful for discussions with S. L. Krug, O. Trottier and A. Khandelwal. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2023-04853. O.A.v.L. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 772834). This research was undertaken thanks in part to funding provided to the University of Toronto’s Acceleration Consortium from the Canada First Research Excellence Fund, grant number: CFREF-2022-00042. O.A.v.L. has received support as the Ed Clark Chair of Advanced Materials and as a Canada CIFAR AI Chair.References
- Ramakrishnan et al. (2014) R. Ramakrishnan, P. O. Dral, M. Rupp, and O. A. von Lilienfeld, Scientific Data 1 (2014).
- Smith et al. (2017) J. S. Smith, O. Isayev, and A. E. Roitberg, Chem. Sci. 8, 3192 (2017).
- Isert et al. (2022) C. Isert, K. Atz, J. Jiménez-Luna, and G. Schneider, Scientific Data 9, 273 (2022).
- Montavon et al. (2013) G. Montavon, M. Rupp, V. Gobre, A. Vazquez-Mayagoitia, K. Hansen, A. Tkatchenko, K.-R. Müller, and O. A. von Lilienfeld, New Journal of Physics 15, 095003 (2013).
- Khan et al. (2024a) D. Khan, A. Benali, S. Y. Kim, G. F. von Rudorff, and O. A. von Lilienfeld, arXiv preprint arXiv:2405.05961 (2024a).
- Li et al. (2023) K. Li, B. DeCost, K. Choudhary, M. Greenwood, and J. Hattrick-Simpers, npj Computational Materials 9, 55 (2023).
- Strubell et al. (2019) E. Strubell, A. Ganesh, and A. McCallum, in Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, edited by A. Korhonen, D. Traum, and L. Màrquez (Association for Computational Linguistics, Florence, Italy, 2019) pp. 3645–3650.
- Thompson et al. (2021) N. C. Thompson, K. Greenewald, K. Lee, and G. F. Manso, IEEE Spectrum 58, 50 (2021).
- Ramakrishnan et al. (2015) R. Ramakrishnan, P. O. Dral, M. Rupp, and O. A. von Lilienfeld, Journal of Chemical Theory and Computation 11, 2087 (2015), pMID: 26574412, https://doi.org/10.1021/acs.jctc.5b00099 .
- Heinen et al. (2024) S. Heinen, D. Khan, G. F. von Rudorff, K. Karandashev, D. J. A. Arrieta, A. J. Price, S. Nandi, A. Bhowmik, K. Hermansson, and O. A. von Lilienfeld, Machine Learning: Science and Technology 5, 025058 (2024).
- Khan et al. (2024b) D. Khan, A. J. A. Price, M. L. Ach, O. Trottier, and O. A. von Lilienfeld, arXiv preprint arXiv:2402.14793 (2024b).
- Khan et al. (2024c) D. Khan, M. L. Ach, and O. A. von Lilienfeld, arXiv preprint arXiv:2404.16942 (2024c).
- Huang and von Lilienfeld (2016) B. Huang and O. A. von Lilienfeld, The Journal of Chemical Physics 145, 161102 (2016), https://doi.org/10.1063/1.4964627 .
- Bigi et al. (2024) F. Bigi, S. N. Pozdnyakov, and M. Ceriotti, The Journal of Chemical Physics 161, 044116 (2024), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/5.0208746/20078941/044116_1_5.0208746.pdf .
- Behler and Parrinello (2007) J. Behler and M. Parrinello, Phys. Rev. Lett. 98, 146401 (2007).
- Behler (2011) J. Behler, The Journal of Chemical Physics 134, 074106 (2011), https://doi.org/10.1063/1.3553717 .
- Braams and Bowman (2009) B. J. Braams and J. M. Bowman, International Reviews in Physical Chemistry 28, 577 (2009), https://doi.org/10.1080/01442350903234923 .
- Rupp et al. (2012a) M. Rupp, A. Tkatchenko, K.-R. Müller, and O. A. von Lilienfeld, Phys. Rev. Lett. 108, 058301 (2012a).
- Bartók et al. (2013) A. P. Bartók, R. Kondor, and G. Csányi, Phys. Rev. B 87, 184115 (2013).
- Drautz (2019) R. Drautz, Phys. Rev. B 99, 014104 (2019).
- Musil et al. (2021) F. Musil, A. Grisafi, A. P. Bartók, C. Ortner, G. Csányi, and M. Ceriotti, Chemical Reviews 121, 9759 (2021), pMID: 34310133, https://doi.org/10.1021/acs.chemrev.1c00021 .
- Tom et al. (2023) G. Tom, R. J. Hickman, A. Zinzuwadia, A. Mohajeri, B. Sanchez-Lengeling, and A. Aspuru-Guzik, Digital Discovery 2, 759 (2023).
- Khan et al. (2023) D. Khan, S. Heinen, and O. A. von Lilienfeld, The Journal of Chemical Physics 159, 034106 (2023), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/5.0152215/18048072/034106_1_5.0152215.pdf .
- Darby et al. (2022) J. P. Darby, J. R. Kermode, and G. Csányi, npj Computational Materials 8 (2022), 10.1038/s41524-022-00847-y.
- Axilrod and Teller (1943) B. M. Axilrod and E. Teller, The Journal of Chemical Physics 11, 299 (1943), https://doi.org/10.1063/1.1723844 .
- Huang and von Lilienfeld (2020a) B. Huang and O. A. von Lilienfeld, Nature Chemistry 12, 945 (2020a).
- Faber et al. (2018) F. A. Faber, A. S. Christensen, B. Huang, and O. A. von Lilienfeld, The Journal of Chemical Physics 148, 241717 (2018), https://doi.org/10.1063/1.5020710 .
- Rupp et al. (2012b) M. Rupp, A. Tkatchenko, K.-R. Müller, and O. A. von Lilienfeld, Phys. Rev. Lett. 108, 058301 (2012b).
- Christensen et al. (2020) A. S. Christensen, L. A. Bratholm, F. A. Faber, and O. Anatole von Lilienfeld, The Journal of Chemical Physics 152, 044107 (2020), https://doi.org/10.1063/1.5126701 .
- Pozdnyakov et al. (2020) S. N. Pozdnyakov, M. J. Willatt, A. P. Bartók, C. Ortner, G. Csányi, and M. Ceriotti, Phys. Rev. Lett. 125, 166001 (2020).
- Parsaeifard et al. (2021) B. Parsaeifard, D. S. De, A. S. Christensen, F. A. Faber, E. Kocer, S. De, J. Behler, O. A. von Lilienfeld, and S. Goedecker, Machine Learning: Science and Technology 2, 015018 (2021).
- Batatia et al. (2022) I. Batatia, D. P. Kovacs, G. Simm, C. Ortner, and G. Csanyi, in Advances in Neural Information Processing Systems, Vol. 35, edited by S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Curran Associates, Inc., 2022) pp. 11423–11436.
- Kovács et al. (2023) D. P. Kovács, I. Batatia, E. S. Arany, and G. Csányi, The Journal of Chemical Physics 159, 044118 (2023), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/5.0155322/18065016/044118_1_5.0155322.pdf .
- Adamo and Barone (1999) C. Adamo and V. Barone, J. Chem. Phys. 110, 6158 (1999).
- Weigend and Ahlrichs (2005) F. Weigend and R. Ahlrichs, Physical Chemistry Chemical Physics 7, 3297 (2005).
- Sun et al. (2018) Q. Sun, T. C. Berkelbach, N. S. Blunt, G. H. Booth, S. Guo, Z. Li, J. Liu, J. D. McClain, E. R. Sayfutyarova, S. Sharma, et al., Wiley Interdisciplinary Reviews: Computational Molecular Science 8, e1340 (2018).
- Huo and Rupp (2022) H. Huo and M. Rupp, Machine Learning: Science and Technology 3, 045017 (2022).
- Moriwaki et al. (2018) H. Moriwaki, Y.-S. Tian, N. Kawashita, and T. Takagi, Journal of Cheminformatics 10 (2018), 10.1186/s13321-018-0258-y.
- Faber et al. (2017) F. A. Faber, L. Hutchison, B. Huang, J. Gilmer, S. S. Schoenholz, G. E. Dahl, O. Vinyals, S. Kearnes, P. F. Riley, and O. A. von Lilienfeld, Journal of Chemical Theory and Computation 13, 5255 (2017), pMID: 28926232, https://doi.org/10.1021/acs.jctc.7b00577 .
- Mazouin et al. (2022) B. Mazouin, A. A. Schöpfer, and O. A. von Lilienfeld, Mater. Adv. 3, 8306 (2022).
- Chen et al. (2023) K. Chen, C. Kunkel, B. Cheng, K. Reuter, and J. T. Margraf, Chemical Science 14, 4913 (2023).
- Llenga and Gryn’ova (2023) S. Llenga and G. Gryn’ova, The Journal of Chemical Physics 158, 214116 (2023), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/5.0151122/17925578/214116_1_5.0151122.pdf .
- Christensen et al. (2019) A. S. Christensen, F. A. Faber, and O. A. von Lilienfeld, The Journal of Chemical physics 150, 064105 (2019).
- Heinen et al. (2022) S. Heinen, G. F. von Rudorff, and O. A. von Lilienfeld, The Journal of Chemical Physics 157, 221102 (2022), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/5.0112856/16554749/221102_1_online.pdf .
- Harris et al. (2020) C. R. Harris, K. J. Millman, S. J. van der Walt, R. Gommers, P. Virtanen, D. Cournapeau, E. Wieser, J. Taylor, S. Berg, N. J. Smith, R. Kern, M. Picus, S. Hoyer, M. H. van Kerkwijk, M. Brett, A. Haldane, J. F. del Río, M. Wiebe, P. Peterson, P. Gérard-Marchant, K. Sheppard, T. Reddy, W. Weckesser, H. Abbasi, C. Gohlke, and T. E. Oliphant, Nature 585, 357 (2020).
- Virtanen et al. (2020) P. Virtanen, R. Gommers, T. E. Oliphant, M. Haberland, T. Reddy, D. Cournapeau, E. Burovski, P. Peterson, W. Weckesser, J. Bright, S. J. van der Walt, M. Brett, J. Wilson, K. J. Millman, N. Mayorov, A. R. J. Nelson, E. Jones, R. Kern, E. Larson, C. J. Carey, İ. Polat, Y. Feng, E. W. Moore, J. VanderPlas, D. Laxalde, J. Perktold, R. Cimrman, I. Henriksen, E. A. Quintero, C. R. Harris, A. M. Archibald, A. H. Ribeiro, F. Pedregosa, P. van Mulbregt, and SciPy 1.0 Contributors, Nature Methods 17, 261 (2020).
- Lam et al. (2015) S. K. Lam, A. Pitrou, and S. Seibert, in Proceedings of the Second Workshop on the LLVM Compiler Infrastructure in HPC, LLVM ’15 (Association for Computing Machinery, New York, NY, USA, 2015).
- Christensen et al. (2017) A. S. Christensen, F. A. Faber, B. Huang, L. A. Bratholm, A. Tkatchenko, K.-R. Müller, and O. A. von Lilienfeld, “Qml: A python toolkit for quantum machine learning,” (2017).
- Himanen et al. (2020) L. Himanen, M. O. J. Jäger, E. V. Morooka, F. Federici Canova, Y. S. Ranawat, D. Z. Gao, P. Rinke, and A. S. Foster, Computer Physics Communications 247, 106949 (2020).
- Laakso et al. (2023) J. Laakso, L. Himanen, H. Homm, E. V. Morooka, M. O. Jäger, M. Todorović, and P. Rinke, The Journal of Chemical Physics 158 (2023).
- Rupp et al. (2012c) M. Rupp, A. Tkatchenko, K.-R. Müller, and O. A. von Lilienfeld, Phys. Rev. Lett. 108, 058301 (2012c).
- Hansen et al. (2015) K. Hansen, F. Biegler, R. Ramakrishnan, W. Pronobis, O. A. von Lilienfeld, K.-R. Müller, and A. Tkatchenko, The Journal of Physical Chemistry Letters 6, 2326 (2015), pMID: 26113956, https://doi.org/10.1021/acs.jpclett.5b00831 .
- Huang and von Lilienfeld (2020b) B. Huang and O. A. von Lilienfeld, Nature chemistry 12, 945 (2020b).
- Huang et al. (2023) B. Huang, O. A. von Lilienfeld, J. T. Krogel, and A. Benali, Journal of Chemical Theory and Computation 19, 1711 (2023).
- Vapnik (1999) V. Vapnik, The nature of statistical learning theory (Springer science & business media, 1999).
- Schütt et al. (2020) K. T. Schütt, S. Chmiela, O. A. Von Lilienfeld, A. Tkatchenko, K. Tsuda, and K.-R. Müller, Lecture Notes in Physics (2020).
V Supplementary Information: Convolutional many body distribution functional representations
V.1 Kernel based methods

### v2

@@ Abstract
Modern machine learning (ML) models of chemical and materials systems with billions of parameters require vast training datasets and considerable computational efforts. Lightweight kernel or decision tree based methods, however, can be rapidly trained, leading to a considerably lower carbon footprint. We introduce generalized convolutional many-body distribution functionals (cMBDF) as highly compute and data efficient atomic representations for accurate kernels that excel in low-data regimes. Generalizing the MBDF framework, cMBDF encodes local chemical environments in a compact fashion using translationally and rotationally invariant functionals of smooth atom centered Gaussian electron density proxy distributions weighted by interaction potentials. The functional values can be efficiently evaluated by expressing them in terms of convolutions which are calculated via fast Fourier transforms and stored on pre-defined grids. In the generalized form each atomic environment is described using a set of functionals uniformly defined by three integers; many-body, derivative, weighting orders. Irrespective of size/composition, cMBDF atomic vectors remain compact and constant in size for a fixed choice of these orders controlling the structural and compositional resolution. While being up to two orders of magnitude more compact than other popular representations, cMBDF is shown to be more accurate for the learning of various quantum properties such as energies, dipole moments, homo-lumo gaps, heat-capacity, polarizability, optimal exact-exchange admixtures and basis-set scaling factors. Applicability for organic and inorganic chemistry is tested as represented by the QM7b, QM9 and VQM24 data sets. Due to its compactness, model training and testing times are reduced from 23 hours to 8 minutes, implying a corresponding reduction in carbon footprint.
@@ I Introduction
Carbon dioxide emissions from training large language models, for instance, are soon expected to match the monthly emissions of New York City 8.
@@ IV Conclusions
In this work we have introduced a systematically improvable class of compact atomic representations for use throughout chemical compound space. Convolutional many-body distribution functionals (cMBDF) encode the chemical environment of an atom through a set of translationally and rotationally invariant functionals of the atomic density. The functional values are efficiently evaluated via fast Fourier transforms using the convolution theorem which can be evaluated and stored on a pre-defined grid. Weighting functions of various types can be incorporated to capture physical interactions (short and long range) of interest efficiently via the atomic density functionals. The atomic feature vector size is invariant to the cut-offs employed, system size and composition leading to a compact and constant size atomic descriptor. cMBDF is shown to outperform other commonly used representations for the learning of a variety of physical properties across chemical space while remaining nearly two orders of magnitude more compact. Due to its computational efficiency and versatility, cMBDF can lead to significantly faster prototyping, training and testing of quantum machine learning models for a variety of tasks across chemical compound space. Future work will include a study of cMBDF gradients and their applicability across geometrical changes for relaxing geometries, identifying transition states 44 and molecular dynamics.
@@ Acknowledgements.
D.K. is thankful for discussions with S. L. Krug and O. Trottier. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2023-04853. O.A.v.L. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 772834). This research was undertaken thanks in part to funding provided to the University of Toronto’s Acceleration Consortium from the Canada First Research Excellence Fund, grant number: CFREF-2022-00042. O.A.v.L. has received support as the Ed Clark Chair of Advanced Materials and as a Canada CIFAR AI Chair.Supplementary Information: Convolutional many body distribution functional representations
V Kernel based methods